OpenStudy (anonymous):

Green light (λ = 536 nm) strikes a single slit at normal incidence. What width slit will produce a central maximum that is 2.80 cm wide on a screen 1.70 m from the slit?

OpenStudy (anonymous):

 \[\sin\theta = \frac{\lambda}{a}\] \[\tan\theta = \frac{max/2}{L}\] Find the angle from the second equation (that's half the central maximum in the numerator) Then use that angle in the first equation and solve for 'a', the slit width.